OpenStudy (anonymous):

Find the derivative dy/dx. y=7t cos (t^4) dt.

OpenStudy (anonymous):

\[\int\limits_{0}^{\sqrt x} 7t \cos(t^4) dt\]

OpenStudy (anonymous):

@satellite73 lol you are my favorite. This is my last one i promise.

OpenStudy (anonymous):

i think i got 7cos t^4 -t(4t^3 sin(t^4)) for the derivative before I put in the sqrt of x

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

it the question to find the derivative of \[\int\limits_{0}^{\sqrt x} 7t \cos(t^4) dt\]? if so , this is way easier than you think

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

did i get it correct above?

OpenStudy (anonymous):

heck no much much easier

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

before we do that, lets do this what is the derivative of \[\int\limits_{0}^{ x} 7t \cos(t^4) dt\]?

OpenStudy (anonymous):

answer: the derivative of the integral is the "integrand" so it is simply \[ 7x \cos(x^4) \]

OpenStudy (anonymous):

oh so then you just plug in the squareroot?

OpenStudy (anonymous):

what did i do? replaced every \(t\) in the integrand by an \(x\) because the integral is a function of \(x\)

OpenStudy (anonymous):

close, but don't forget the "chain rule"

OpenStudy (anonymous):

yo do not have \[\int\limits_{0}^{x} 7t \cos(t^4) dt\] you have \[\int\limits_{0}^{\sqrt x} 7t \cos(t^4) dt\]which is like a composite function \[F(\sqrt{x})\]

OpenStudy (anonymous):

so to take the derivative, replace every \(t\) by \(\sqrt{x}\) and then don't forget to multiply the result by the derivative of \(\sqrt{x}\) namely \(\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

so it is 7/2 cos x^2

OpenStudy (anonymous):

lets see i didn't do it

OpenStudy (anonymous):

oh sorry but thank you!

OpenStudy (anonymous):

\[7\sqrt{x}\cos(\sqrt{x}^4)\times \frac{1}{2\sqrt{x}}\] \[\frac{7}{2}\cos(x^2)\]

OpenStudy (anonymous):

thank you it totally makes sense now!

OpenStudy (anonymous):

yup, you got it much easier than you thought i bet

OpenStudy (anonymous):

i thougth about it wayy to hard.

OpenStudy (anonymous):

it is math, don't think too hard

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

yw